OpenStudy (anonymous):

Find t(1) for this arithmetic sequence.

OpenStudy (anonymous):

\[d=6, S _{n}=574 , n=14\]

OpenStudy (anonymous):

Sorry, for each arithmetic SERIES.

OpenStudy (cwrw238):

sum of n terms = (n/2)[2a + (n-1)d] where a = t(1) 574 = 7[2a + 13*6] solve for a

OpenStudy (anonymous):

That makes sense, but the equation that I was given in class was: \[S{n}=\frac{ n }{ 2 } (t{1} + t{n})\] Is that the same thing? :S

OpenStudy (cwrw238):

tn is nth term which is a + (n-1)d - yes

OpenStudy (anonymous):

Okay, thank you so much :)

OpenStudy (cwrw238):

yw tn is last term